Use natural logarithms to solve the equation. Round to the nearest thousandth.
5e^(2x + 11)=30

Answer by lwsshak3(11628)

Use natural logarithms to solve the equation. Round to the nearest thousandth.
5e^(2x + 11)=30
divide by 5
e^(2x + 11)=6
(2x+11)lne=ln6
lne=1
2x+11=ln6
2x=ln6-11
x=(ln6-11)/2
x≈-4.604